Use the geometric sequence of numbers 1, 1/2, 1/4, 1/8,…to find the following:
a) What is r, the ratio between 2 consecutive terms?
Answer:
Show work in this space.



b) Using the formula for the sum of the first n terms of a geometric series, what is the sum of the first 10 terms? Please round your answer to 4 decimals.
Answer:
Show work in this space.


c) Using the formula for the sum of the first n terms of a geometric series, what is the sum of the first 12 terms? Please round your answer to 4 decimals.
Answer:
Show work in this space.



d) What observation can make about these sums? In particular, what number does it appear that the sum will always be smaller than?
Answer:

Found 2 solutions by AnlytcPhil, funmath:
Answer by AnlytcPhil(1806)

Use the geometric sequence of numbers 1, 1/2, 1/4, 1/8,…to find 
the following:
a) What is r, the ratio between 2 consecutive terms? 
Answer: 
Show work in this space. 

2nd term divided by 1st term = 1/2 ÷ 1 = 1/2
3rd term divided by 2nd term = 1/4 ÷ 1/2 = 1/4 × 2/1 = 2/4 = 1/2
4th term divided by 3rd term = 1/8 ÷ 1/4 = 1/8 × 4/1 = 4/8 = 1/2

Those all have to be equal the same for it to be a geometric 
sequence. They are so it is.  That common ratio is therefore the 
number r = 1/2

================================================================

b) Using the formula for the sum of the first n terms of a 
geometric series, what is the sum of the first 10 terms? Please 
round your answer to 4 decimals.
Answer: 
Show work in this space. 

the sum of the first n terms is given by the formula

      a1(1 - rn)
Sn = ------------  
        1 - r

where a1 = the first term or 1. r = 1/2, and n = 10

       1[1 - (1/2)10]
S10 = -----------------  
          1 - 1/2


         1 - (1/2)10
S10 = -----------------  
             1/2

         1 - 1/210
S10 = -----------------  
             1/2

Multiply top and bottom by 2

        2(1 - 1/210)
S10 = ----------------  
           2(1/2)

        2 - 2/210
S10 = ----------------  
            1

S10 = 2 - 1/29

S10 = 2 - 1/512

S10 = 1024/2 = 1/512

S10 = 1023/512 = 1.998046875, or 1.9980 to 
4 decimals

================================================================

c) Using the formula for the sum of the first 
n terms of a geometric series, what is the sum 
of the first 12 terms? Please round your answer 
to 4 decimals.
Answer: 
Show work in this space. 

      a1(1 - rn)
Sn = ------------  
        1 - r

where a1 = the first term or 1. r = 1/2, and 
this time n = 12

       1[1 - (1/2)12]
S12 = -----------------  
          1 - 1/2


         1 - (1/2)12
S12 = -----------------  
             1/2

         1 - 1/212
S12 = -----------------  
             1/2

Multiply top and bottom by 2

        2(1 - 1/212)
S12 = ----------------  
           2(1/2)

        2 - 2/212
S12 = ----------------  
            1

S12 = 2 - 1/211

S12 = 2 - 1/2048

S12 = 4096/2 = 1/2048

S12 = 4095/2048 = 1.9999511719, or 1.9999 to 
4 decimals

================================================================

d) What observation can make about these sums? 
In particular, what number does it appear that 
the sum will always be smaller than?
Answer:

1.9980 and 1.9999 are very close to 2 

If we just leave the n without replacing it in 
the formula,

      a1(1 - rn)
Sn = ------------  
        1 - r

where a1 = the first term or 1. r = 1/2, and 
n = 10

       1[1 - (1/2)n]
Sn = -----------------  
          1 - 1/2

        1 - (1/2)n
Sn = -----------------  
           1/2

        1 - 1/2n
Sn = -----------------  
          1/2

Multiply top and bottom by 2

      2(1 - 1/2n)
Sn = -------------  
         2(1/2)

      2 - 2/2n
Sn = ----------  
          1

Sn = 2 - 1/2n-1

The fraction 1/2n-1 is subtracted from 2 in the
formula.  It gets smaller and smaller as n gets
larger because its denominator 2n-1 gets larger 
and larger. So we are each time subtracting less 
and less from 2. But the fraction will never 
reach 0, so Sn will always be just a tiny 
fraction under 2

Edwin


Answer by funmath(2933)

a) a(sub2)/a(sub1)=a(sub3)/a(sub2)=...=r
a(sub1)=1;a(sub2)=1/2
r=(1/2)/1=1/2
-------------------------------------
b) S(subn)=a(sub1)((1-r^n)/(1-r))
n=10;a(sub1)=1;r=1/2
S(sub10)=(1)((1-(1/2)^10)/(1-1/2))
=(1-1/1024)/(1-1/2)
=(1023/1024)/(1/2)
=(1023/1024)*(2/1)
=1023/512
=1.9980
----------------------------------------
c)S(subn)=a(sub1)((1-r^n)/(1-r))
n=12;a(sub1)=1;r=1/2
S(sub12)=(1)(1-(1/2)^12)/(1-1/2)
=(1-1/4096)/(1-1/2)
=(4095/4096)/(1/2)
=(4095/4096)*(2/1)
=4095/2048
=1.9995
-----------------------------------------
d) It's getting closer to 2, but not equalling 2. It appears that it will always be less than 2.